Please help with this problem:
Circle one: f(x)=6x^2-36x+59 has {zero, one or two} roots.
Can you please show me how you got the answer
Thank you.

Answer by jim_thompson5910(35256)

From f(x)=6x^2-36x+59, we see that a = 6, b = -36 and c = 59

This is because the general form is f(x) = ax^2 + bx + c

-------------------------------------------------------

Use the discriminant formula D = b^2 - 4ac


D = b^2 - 4ac

D = (-36)^2 - 4(6)(59) ... plug in a = 6, b = -36 and c = 59

D = -120


The discriminant is negative, so there are zero real solutions.